OpenStudy (anonymous):

The legs of a right triangle are 8 and 9 feet long, respectively. What is the length of the hypotenuse? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

do u know the pythagorean theorem? it is used in all right triangles and relates the sides to the hypotenuse...

OpenStudy (anonymous):

@dpaInc So 8^2=64, 9^2=81. 64+81=145. Do I have to find the square root?

OpenStudy (anonymous):

yes... because that 145 is the hypotenuse squared from \(\large a^2+b^2=hypotenuse^2 \)

OpenStudy (anonymous):

So then it would be 12.04 correct?

OpenStudy (anonymous):

that looks good....

OpenStudy (anonymous):

thanks much appreciated!

OpenStudy (anonymous):

yw... :)